kacchanforlife:

If you have points (3,6) and (2,1) on a graph, what is the equation of the line in standard form? 7x+y=13 y-8=7(x-3) 7x-y=13 If you have points (1,4) and (-1,2) on a graph, what is the equation of the line in slope intercept form?

kacchanforlife:

idfkaa:

@oliver69

kacchanforlife:

@idfkaa wrote:
@oliver69
he's not online

issy13:

(y - 6) = 1(x - 3).

kacchanforlife:

for which one?

issy13:

for ur question?

kacchanforlife:

that isnt an answer choice

issy13:

oh than idk

issy13:

1 attachment
kacchanforlife:

?

OLIVER69:

For the first image, it would be x+y=56. This and the first one were the only ones in standard form. And as for why this is correct because there is a 0 in the x row, which is above the 56.

OLIVER69:

As for the second, it would be y=5x+5 because it has a slops of 5 and the y-intercept is 5. I hope I could help.

OLIVER69:

As for this question, " If you have points (3,6) and (2,1) on a graph, what is the equation of the line in standard form?" Your answer would be y-8=7(x-3)